OpenStudy (anonymous):

Which of the following points lies on the circle whose center is at the origin and whose radius is 13?

OpenStudy (anonymous):

x^2+y^2=169 let's check all 25+144=169 (a) is 13+13=26 (b) is NOT 36+49=85 (b) is NOT only (a)